# Constrained Stochastic Recursive Momentum Successive Convex Approximation

### v2

@@ Abstract
We consider stochastic optimization problems with functional constraints, such as those arising in trajectory generation, sparse approximation, and robust classification. To this end, we put forth a recursive momentum-based accelerated successive convex approximation (SCA) algorithm. At each iteration, the proposed algorithm entails constructing convex surrogates of the stochastic objective and the constraint functions, and solving the resulting convex optimization problem. A recursive update rule is employed to track the gradient of the stochastic objective function, which contributes to variance reduction and hence accelerates the algorithm convergence. A key ingredient of the proof is a new parameterized version of the standard Mangasarian-Fromowitz Constraints Qualification, that allows us to bound the dual variables and hence obtain problem-dependent bounds on the rate at which the iterates approach an $\epsilon$-stationary point. Remarkably, the proposed algorithm achieves near-optimal stochastic first order (SFO) complexity, almost at par with that achieved by state-of-the-art stochastic optimization algorithms for solving unconstrained problems. As an example, we detail a obstacle-avoiding trajectory optimization problem that can be solved using the proposed algorithm and show that its performance is superior to that of the existing algorithms used for trajectory optimization. The performance of the proposed algorithm is also shown to be comparable to that of a specialized sparse classification algorithm applied to a binary classification problem.
@@ I-A Background
Existing approaches to solve ($\mathcal{P}$) include the primal-dual method [9], penalty method [10], proximal point methods [11, 12], augmented Lagrangian method [13], and several Successive Convex Approximation (SCA) variants [14, 15, 16, 17, 18]. The performance of these algorithms is usually measured in terms of their oracle complexity, which is the number of calls to the stochastic first-order (SFO) oracle required to reach a point that satisfies the Karush-Kuhn Tucker (KKT) conditions to within a tolerance of $\epsilon$. The primal-dual method proposed in [9] iteratively updates both primal and dual variables based on a stochastic approximation to augmented Lagrangian function, achieving an oracle complexity of $\mathcal{O}(\epsilon^{-2.5})$. A rate of $\mathcal{O}(\epsilon^{-3.5})$ is achieved by [10] which utilizes a penalty functional to reformulate ($\mathcal{P}$) as an unconstrained optimization problem. An improved rate of $\mathcal{O}(\epsilon^{-2})$ is achieved by the Inexact Constrained Proximal Point Algorithm (ICPPA) [12] and level-constrained stochastic proximal gradient (LCSPG) methods [11], both of which solve a sequence of strongly convex subproblems, obtained by adding a quadratic term to the objective and constraint functions. Even though [12, 11] achieve $\mathcal{O}(\epsilon^{-2})$ under a standard Lipschitz smoothness assumption on $U$, the augmented Lagrangian method [13] achieves the same rate under milder mean-sqaured smoothness (MSS) assumption. A related level-constrained proximal point (LCPP) method [8] also achieved the same rate but was customized to only handle sparsity inducing constraint functions.
@@
Different from these approaches, SCA represents a more flexible framework for solving generic non-convex constrained optimization problems. It is an iterative approach, where at every iteration, the non-convex objective and constraint functions are replaced with their convex surrogates, tailored to the problem at hand [19, 20]. Stochastic SCA variants have been proposed in [21, 14, 15, 16, 17, 18] for solving ($\mathcal{P}$) as well as its more general version with stochastic constraints. These works however only report asymptotic convergence results, and do not provide a comprehensive oracle complexity analysis. Non-asymptotic rates are only known for the classical SCA algorithm when solving ($\mathcal{P}$) with only convex constraints [22, 23, 24, 25], where the latter work achieved the state-of-the-art rate of $\mathcal{O}(\epsilon^{-2})$.
@@
We remark that while the oracle complexity of ${\mathcal{O}}(\epsilon^{-2})$ achieved by various algorithms that solve ($\mathcal{P}$) may be optimum under the Lipschitz smoothness assumption on $U$, it can be improved to $\mathcal{O}(\epsilon^{-3/2})$ under the MSS assumption on $f$ [26, (4)]. Indeed, using momentum or variance reduction techniques [27, 28, 29, 30, 31], it is well-known that the optimal oracle complexity of $\mathcal{O}(\epsilon^{-3/2})$ can be achieved for unconstrained or convex-constrained version of ($\mathcal{P}$). Interestingly, variance reduction is applied to the level-constrained proximal algorithm in [11] for finite-sum problems but the rate achieved is only ${\mathcal{O}}(\epsilon^{-2})$, with a remark that it can be improved to $\mathcal{O}(\epsilon^{-3/2})$ if $f({\mathbf{x}},{\boldsymbol{\xi}})$ is Lipschitz smooth, but without any proof [11, Remark 5].
@@
Before concluding this section, we remark that analysis of optimization problems with non-convex constraints inevitably requires some assumptions called the constraint qualifications (CQ). These CQs are essential regularity conditions that ensure that the constraint functions are well-behaved, so that the intermediate updates or subproblems are well-defined. Technically, these CQs are utilized to ensure that the intermediate dual variables remain bounded across iterations. Commonly used CQs include the Non-singularity condition (NSC) [9], Linear Independence CQ (LICQ) [32], Linear Independence regularity condition (LIRC) [18], and most commonly the Mangasarian-Fromovitz CQ (MFCQ) condition [12, 11, 8]. It was observed in [11, Sec. 3.1] that the classical MFCQ only implies the existance of a bound $B$ on the dual variables, but does not relate $B$ to the other problem parameters, leaving the possibility of $B$ getting arbitrarily large. It was suggested in [11] to use a stronger CQ such as the strong feasibility CQ (SFCQ) so as to ensure a problem parameter-dependent oracle complexity bound. In this work, we instead utilize a stronger parameterized version of MFCQ to the same effect, i.e., bound the dual variables in terms of problem-dependent quantities.
@@ I-B Contributions
  • Accelerated Stochastic SCA algorithm: We introduce the STOchastic Recursive Momentum (STORM) updates from [33, 27] into the SCA framework to achieve near-optimal SFO complexity. Drawing inspiration from these works, the proposed Constrained STORM SCA (CoSTA) algorithm incorporates momentum-enhanced updates with adaptive step size.
@@
Smoothness Remarks
@@
${{\tilde{\mathcal{O}}}(\epsilon^{-3/2})}$ Parameterized MFCQ $f$ is Smooth SCA-based
@@
MLALM[13] $\mathcal{O}(\epsilon^{-2})$ extended MFCQ $f$ is MSS Augmented Lagrangian
@@ II Motivating Examples
$\left\|{\mathbf{x}}_{i}(\tau+1)-{\mathbf{x}}_{i}(\tau)-\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\Delta t\right\|$ $\leq\left\|{\mathbf{x}}_{i}(\tau+1)-{\mathbf{x}}_{i}(\tau)-{\mathbb{E}}\left[\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\right]\Delta t\right\|$ (2)
@@ III-B Proposed Algorithm
We now described the proposed successive convex optimization algorithm in detail. The algorithm is initialized at an arbitrary feasible point ${\mathbf{x}}_{1}={\mathbf{x}}_{0}\in{\mathcal{K}}$ that satisfies ${\mathbf{g}}({\mathbf{x}}_{1})\leq 0$. At iteration $t$ and given iterate ${\mathbf{x}}_{t}$, we construct a strongly convex and smooth surrogate ${\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$ of the objective function $f({\mathbf{x}},{\boldsymbol{\xi}}_{t})$ as well as smooth and convex surrogates $\{{\tilde{g}}_{j}({\mathbf{x}},{\mathbf{x}}_{t})\}_{j=1}^{J}$ (collected into the vector-valued function ${\tilde{{\mathbf{g}}}}({\mathbf{x}},{\mathbf{x}}_{t})$) of the constraint functions $\{g_{j}({\mathbf{x}})\}_{j=1}^{J}$. These surrogates are subsequently used to formulate the following convex sub-problem that must be solved at every iteration:
@@
${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t}):={\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$ (6)
$+(1-\beta_{t})\langle{\mathbf{x}}-{\mathbf{x}}_{t},{\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t})\rangle$
@@ III-C1 Assumptions on ($\mathcal{P}$)
Application of CoSTA requires ($\mathcal{P}$) to satisfy five key properties: problem should be well-defined on ${\mathcal{X}}$, initial ${\mathbf{x}}_{1}$ should be feasible, $f$ and $h_{i}$s should be smooth and Lipschitz-continuous, variance of the objective gradient should be bounded, and finally MFCQ.
@@
and functions $\{h_{i}\}_{i=1}^{I}$ are $L$-smooth and $G$-Lipschitz, i.e.,
@@
A4.
@@
A5.
@@
Assumptions A1 and A2 are standard and introduced to simplify the analysis. Implicit within A2 is the requirement that $F^{\star}$ is bounded from below. The smoothness requirement in Assumption A3 is again standard in the context of gradient-based algorithms. Note that Assumption A3 implies that $f$ is MSS and that $U$ is $L$-smooth. The $G$-Lipschitz condition in Assumption A3 implies that $G_{t}\leq G$ for all $t\geq 1$. Assumption A4 is standard in the context of proximal stochastic gradient descent. A more relaxed version, where the gradient boundedness is only required at a specific point rather than at all points in ${\mathcal{X}}$ has also been used in the literature [44], but its applicability to the non-convex constrained case remains an open problem. The MFCQ assumption in A5 is critical and also common in the context of constrained optimization. The condition stems from the need to solve the convex optimization problem ($\mathcal{P}_{t}$) at every iteration. The subsequent analysis, as well as the analysis of most widely used convex optimization solvers, necessitates that Slater’s condition be satisfied for the convex subproblem, primarily to ensure that the dual variables stay bounded. The MFCQ assumption on ($\mathcal{P}$) is required to ensure that Slater’s condition holds for ($\mathcal{P}_{t}$).
In the literature, the classical MFCQ condition, which corresponds to $\chi=0$ and $\rho\rightarrow 0$, is commonly encountered, but is known to yield asymptotic convergence results, as in [19]. This is because the classical MFCQ only implies that the dual variables associated with ($\mathcal{P}_{t}$) are bounded by an unknown parameter that may potentially depend on the solution path, as was also observed in [12, Remark 5]. However, the parameterized version of MFCQ in Assumption A5 ensures that this bound is computable and depends solely on the initial conditions and the specific problem parameters such as $L$ and $\rho$. Intuitively, the vector ${\mathbf{d}}_{{\tilde{\mathbf{x}}}}$ can be interpreted as a direction along which one could move a sufficiently small distance, starting from ${\tilde{\mathbf{x}}}$, and stay feasible. We will formalize this intuition in Sec. IV and establish that Assumption A5 implies the existence of a Slater point for ($\mathcal{P}_{t}$). The existence of such a Slater point immediately yields a bound on the dual variables.

### v3

@@ Abstract
We consider stochastic optimization problems with non-convex functional constraints, such as those arising in trajectory generation, sparse approximation, and robust classification. To this end, we put forth a recursive momentum-based accelerated successive convex approximation (SCA) algorithm. At each iteration, the proposed algorithm entails constructing convex surrogates of the stochastic objective and the constraint functions, and solving the resulting convex optimization problem. A recursive update rule is employed to track the gradient of the stochastic objective function, which contributes to variance reduction and hence accelerates the algorithm convergence. A key ingredient of the proof is a new parameterized version of the standard Mangasarian-Fromowitz Constraints Qualification, that allows us to bound the dual variables and hence obtain problem-dependent bounds on the rate at which the iterates approach an $\epsilon$-stationary point. Remarkably, the proposed algorithm achieves near-optimal stochastic first-order (SFO) complexity with adaptive step sizes and optimal SFO complexity with non-adaptive step sizes at par with that achieved by state-of-the-art stochastic optimization algorithms for solving unconstrained problems. As an example, we detail a obstacle-avoiding trajectory optimization problem that can be solved using the proposed algorithm and show that its performance is superior to that of the existing algorithms used for trajectory optimization. The performance of the proposed algorithm is also shown to be comparable to that of a specialized sparse classification algorithm applied to a binary classification problem.
@@ I-A Background
Existing approaches to solve ($\mathcal{P}$) include the primal-dual method [10], penalty method [11], proximal point methods [12, 13], augmented Lagrangian method [14], and several Successive Convex Approximation (SCA) variants [15, 16, 17, 18, 19]. The performance of these algorithms is usually measured in terms of their oracle complexity, which is the number of calls to the stochastic first-order (SFO) oracle required to reach a point that satisfies the Karush-Kuhn Tucker (KKT) conditions to within a tolerance of $\epsilon$. The primal-dual method proposed in [10] iteratively updates both primal and dual variables based on a stochastic approximation to augmented Lagrangian function, achieving an oracle complexity of $\mathcal{O}(\epsilon^{-2.5})$. A rate of $\mathcal{O}(\epsilon^{-3.5})$ is achieved by [11] which utilizes a penalty functional to reformulate ($\mathcal{P}$) as an unconstrained optimization problem. An improved rate of $\mathcal{O}(\epsilon^{-2})$ is achieved by the Inexact Constrained Proximal Point Algorithm (ICPPA) [13] and level-constrained stochastic proximal gradient (LCSPG) methods [12], both of which solve a sequence of strongly convex subproblems, obtained by adding a quadratic term to the objective and constraint functions. Both [13] and [12] achieve $\mathcal{O}(\epsilon^{-2})$ SFO complexity using fixed $\epsilon$-dependent step-sizes under a standard Lipschitz smoothness assumption on $U$. A related level-constrained proximal point (LCPP) method [9] also achieved the a rate of ${\mathcal{O}}(\epsilon^{-2})$ but was customized to only handle sparsity inducing constraint functions. Improving on these works, the augmented Lagrangian method (MLALM) [14] requires a weaker mean-squared smoothness (MSS) assumption but still achieves the optimal rate of $\mathcal{O}(\epsilon^{-3/2})$, albeit for a fixed $\epsilon$-dependent step-size. In contrast, modern unconstrained stochastic optimization methods such as STORM and its variants have moved to using adaptive step-sizes that are generally easier to tune [20].
@@
Different from these approaches, SCA represents a more flexible framework for solving generic non-convex constrained optimization problems. It is an iterative approach, where at every iteration, the non-convex objective and constraint functions are replaced with their convex surrogates, tailored to the problem at hand [21, 22]. Stochastic SCA variants have been proposed in [23, 15, 16, 17, 18, 19] for solving ($\mathcal{P}$) as well as its more general version with stochastic constraints. These works however only report asymptotic convergence results, and do not provide a comprehensive oracle complexity analysis. Non-asymptotic rates are only known for the classical SCA algorithm when solving ($\mathcal{P}$) with only convex constraints [24, 25, 26, 27], [28], where the works [27] [28] achieved the state-of-the-art rate of $\mathcal{O}(\epsilon^{-2})$.
@@
We remark that while the oracle complexity of ${\mathcal{O}}(\epsilon^{-2})$ achieved by various algorithms that solve ($\mathcal{P}$) may be optimum under the Lipschitz smoothness assumption on $U$, it can be improved to $\mathcal{O}(\epsilon^{-3/2})$ under the MSS assumption on $f$ [29, (4)]. Indeed, using momentum or variance reduction techniques [30, 31, 32, 33, 34], it is well-known that the optimal oracle complexity of $\mathcal{O}(\epsilon^{-3/2})$ can be achieved for unconstrained or convex-constrained version of ($\mathcal{P}$). Interestingly, variance reduction is applied to the level-constrained proximal algorithm in [12] for finite-sum problems but the rate achieved is only ${\mathcal{O}}(\epsilon^{-2})$, with a remark that it can be improved to $\mathcal{O}(\epsilon^{-3/2})$ if $f({\mathbf{x}},{\boldsymbol{\xi}})$ is Lipschitz smooth in ${\mathbf{x}}$ for all ${\boldsymbol{\xi}}$, but without any proof [12, Remark 5]. As noted earlier, only the recently published work [14] achieves the optimal rate of $\mathcal{O}(\epsilon^{-3/2})$, although under a restrictive condition requiring a fixed $\epsilon$-dependent step size.
@@
Before concluding this section, we remark that analysis of optimization problems with non-convex constraints inevitably requires some assumptions called the constraint qualifications (CQ). These CQs are essential regularity conditions that ensure that the constraint functions are well-behaved, so that the intermediate updates or subproblems are well-defined. Technically, these CQs are utilized to ensure that the intermediate dual variables remain bounded across iterations. Commonly used CQs include the Non-singularity condition (NSC) [10], Linear Independence CQ (LICQ) [35], Linear Independence regularity condition (LIRC) [19], and most commonly the Mangasarian-Fromovitz CQ (MFCQ) condition [13, 12, 9]. It was observed in [12, Sec. 3.1] that the classical MFCQ only implies the existence of a bound $B$ on the dual variables, but does not relate $B$ to the other problem parameters, leaving the possibility of $B$ getting arbitrarily large. It was suggested in [12] to use a stronger CQ such as the strong feasibility CQ (SFCQ) so as to ensure a problem parameter-dependent oracle complexity bound. An extended version of MFCQ is used in [14] in order to establish that the Lagrange multiplier stays bounded for all iterations. However, the extended MFCQ does not yield an explicit bound on the Lagrange multipliers, and hence the final SFO complexity in [14] depends on an unknown (but finite) quantity $\tilde{\Lambda}$. In contrast, the slightly stronger parameterized MFCQ introduced here allows us to provide explicit bounds on the dual variables in terms of problem-dependent quantities using simpler proofs.
@@ I-B Contributions
  • Accelerated Stochastic SCA algorithm: we incorporate the STOchastic Recursive Momentum (STORM) updates from [20, 30] into the SCA framework to achieve near-optimal SFO complexity under adaptive step sizes and optimal SFO complexity under non-adaptive step sizes. Drawing inspiration from these works, the proposed Constrained STORM SCA (CoSTA) algorithm leverages momentum-enhanced updates to improve convergence rates.
@@
Smoothness
Adaptive
stepsize
Remarks
@@
${\tilde{\mathcal{O}}}(\epsilon^{-3/2})$
${\mathcal{O}}(\epsilon^{-3/2})$
Parameterized MFCQ
$f$ is Smooth
$f$ is MSS
SCA-based, ${\tilde{\mathcal{O}}}(\epsilon^{-3/2})$ with adaptive step-size
and ${\mathcal{O}}(\epsilon^{-3/2})$ with non-adaptive step-size
@@
MLALM[14] $\mathcal{O}(\epsilon^{-3/2})$ Extended MFCQ $f$ is MSS
${\mathcal{O}}(\epsilon^{-3/2})$ but depending
on an unknown finite $\tilde{\Lambda}$
@@ II Motivating Examples
$\left\|{\mathbf{x}}_{i}(\tau+1)-{\mathbf{x}}_{i}(\tau)-\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\Delta t\right\|$ (2)
$\leq\left\|{\mathbf{x}}_{i}(\tau+1)-{\mathbf{x}}_{i}(\tau)-{\mathbb{E}}\left[\vartheta({\mathbf{x}}_{i}(\tau),{\boldsymbol{\xi}})\right]\Delta t\right\|$
@@ III-B Proposed Algorithm
We now describe the proposed successive convex optimization algorithm in detail. The algorithm is initialized at an arbitrary feasible point ${\mathbf{x}}_{1}={\mathbf{x}}_{0}\in{\mathcal{K}}$ that satisfies ${\mathbf{g}}({\mathbf{x}}_{1})\leq 0$. At iteration $t$ and given iterate ${\mathbf{x}}_{t}$, we construct a strongly convex and smooth surrogate ${\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$ of the objective function $f({\mathbf{x}},{\boldsymbol{\xi}}_{t})$ as well as smooth and convex surrogates $\{{\tilde{g}}_{j}({\mathbf{x}},{\mathbf{x}}_{t})\}_{j=1}^{J}$ (collected into the vector-valued function ${\tilde{{\mathbf{g}}}}({\mathbf{x}},{\mathbf{x}}_{t})$) of the constraint functions $\{g_{j}({\mathbf{x}})\}_{j=1}^{J}$. These surrogates are subsequently used to formulate the following convex sub-problem that must be solved at every iteration:
@@
${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t}):={\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+(1-\beta_{t})\langle{\mathbf{x}}-{\mathbf{x}}_{t},{\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t})\rangle$ (6)
@@ III-C1 Assumptions on ($\mathcal{P}$)
Application of CoSTA requires ($\mathcal{P}$) to satisfy five key properties: problem should be well-defined on ${\mathcal{X}}$, initial ${\mathbf{x}}_{1}$ should be feasible, $f,g_{j}$s and $h_{i}$s should be smooth and Lipschitz-continuous, variance of the objective gradient should be bounded, and finally MFCQ.
@@
for all ${\boldsymbol{\xi}}$.
A4.
Functions $\{h_{i}\}_{i=1}^{I}$ are $L$-smooth and $G$-Lipschitz, i.e.,
@@
and functions $\{g_{j}\}_{j=1}^{J}$ are $L$-smooth and $G$-Lipschitz,
$\left\|\nabla g_{j}({\mathbf{x}})-\nabla g_{j}({\mathbf{y}})\right\|$ $\leq L\left\|{\mathbf{x}}-{\mathbf{y}}\right\|,$ (16)
$\lvert g_{j}({\mathbf{x}})-g_{j}({\mathbf{y}})\rvert$ $\leq G\left\|{\mathbf{x}}-{\mathbf{y}}\right\|$ (17)
A5.
@@
A6.
Mean Squared Smooth (MSS):
${\mathbb{E}}[\left\|\nabla f({\mathbf{x}},{\boldsymbol{\xi}})-\nabla f({\mathbf{y}},{\boldsymbol{\xi}})\right\|^{2}]$ $\leq L^{2}\left\|{\mathbf{x}}-{\mathbf{y}}\right\|^{2}$ (18)
for all ${\boldsymbol{\xi}}$ and Mean Squared Lipschitz (MSL)
${\mathbb{E}}[\lvert f({\mathbf{x}},{\boldsymbol{\xi}})-f({\mathbf{y}},{\boldsymbol{\xi}})\rvert]$ $\leq G\left\|{\mathbf{x}}-{\mathbf{y}}\right\|$ (19)
for all ${\boldsymbol{\xi}}$.
A7.
@@
Assumptions A1 and A2 are standard and introduced to simplify the analysis. Implicit within A2 is the requirement that $F^{\star}$ is bounded from below. The smoothness requirement in Assumption A3 and A4 are again standard in the context of gradient-based algorithms. Note that Assumption A3 implies that $f$ is MSS and $U$ is $L$-smooth. The $G$-Lipschitz condition in Assumption A3 implies that $G_{t}\leq G$ for all $t\geq 1$. Note that the MSS and MSL conditions in Assumption A6 are milder than their counterparts in Assumption A3. Assumption A5 is standard in the context of proximal stochastic gradient descent. A more relaxed version, where the gradient boundedness is only required at a specific point rather than at all points in ${\mathcal{X}}$ has also been used in the literature [46], but its applicability to the non-convex constrained case remains an open problem. The MFCQ assumption in A7 is critical and also common in the context of constrained optimization. The condition stems from the need to solve the convex optimization problem ($\mathcal{P}_{t}$) at every iteration. The subsequent analysis, as well as the analysis of most widely used convex optimization solvers, necessitates that Slater’s condition be satisfied for the convex subproblem, primarily to ensure that the dual variables stay bounded. The MFCQ assumption on ($\mathcal{P}$) is required to ensure that Slater’s condition holds for ($\mathcal{P}_{t}$).
In the literature, the classical MFCQ condition, which corresponds to $\chi=0$ and $\rho\rightarrow 0$, is commonly encountered, but is known to yield asymptotic convergence results, as in [21]. This is because the classical MFCQ only implies that the dual variables associated with ($\mathcal{P}_{t}$) are bounded by an unknown parameter that may potentially depend on the solution path, as was also observed in [13, Remark 5]. Another approach, as also mentioned in Sec. I-A, is the extended MFCQ used in [14] and corresponds to using $\chi=0$ and $\rho>0$. While the extended MFCQ guarantees a bound on the dual variables, it does not yield an explicit expression for the bound, which leaves the final SFO complexity depending on an unknown quantity $\tilde{\Lambda}$ in [14]. In contrast, the parameterized version of MFCQ in Assumption A7 ensures that this bound is computable and depends solely on the initial conditions and the specific problem parameters such as $L$ and $\rho$. Intuitively, the vector ${\mathbf{d}}_{{\tilde{\mathbf{x}}}}$ can be interpreted as a direction along which one could move a sufficiently small distance, starting from ${\tilde{\mathbf{x}}}$, and stay feasible. We will formalize this intuition in Sec. IV and establish that Assumption A7 implies the existence of a Slater point for ($\mathcal{P}_{t}$). The existence of such a Slater point immediately yields a bound on the dual variables.